# thanks: These authors contributed equallythanks: These authors contributed equally

### v1

@@ 3.4.2 Variant 2: Joint parity measurements
On closer inspection of Table 1, examining the outcomes for $\sigma_{a},\sigma_{b}$, one may see that there is indeed a redundancy that can be eliminated. We see that for the four mode $c$ states, the product $\sigma_{a}\sigma_{b}$ yields the same value. This suggests that by performing a joint parity measurement of $\sigma_{ab}=\sigma_{a}\sigma_{b}$, without measuring $\sigma_{a},\sigma_{b}$ individually, one can achieve the same result as Table 1. Such joint parity measurements have also been demonstrated experimentally [43, 47]. Such a method can be calculated by applying the joint parity projection operator $P_{\sigma_{ab}}$ on the state (20).
@@
Similarly to (26), this involves a superposition of two versions of the initial state. The difference here is that modes $a$ and $b$ are left in one of the Bell states. The superposition of these terms can be further collapsed by distinguishing between the amplitudes $\chi$ and $\bar{\chi}$, using a procedure similar to that described in Sec. 3.4.1. In Table 2, we summarize the various outcomes for the joint parity measurement and the displaced measurement.
@@
$\mu|0_{L}\rangle_{\beta}+\nu|1_{L}\rangle_{\beta}$ $+$ $+$ $I_{L}$
$\mu|0_{L}\rangle_{\beta}-\nu|1_{L}\rangle_{\beta}$ $+$ -- $Z_{L}$
$\mu|1_{L}\rangle_{\beta}+\nu|0_{L}\rangle_{\beta}$- $+$ $X_{L}$
$-\mu|1_{L}\rangle_{\beta}+\nu|0_{L}\rangle_{\beta}$- -- $X_{L}Z_{L}$
@@ 3.5 Correction operation
4 Teleportation performance using single-mode parity measurements
In this section, we numerically analyze the teleportation protocol as introduced in Sec. 3. Of the two variants of the teleportation protocol involving the single-mode parity measurements (Variant 1, Sec. 3.4.1) and the joint parity measurements (Variant 2, Sec. 3.4.2), in this section we consider the former. The latter will be analyzed in Sec. 5.
We have numerically verified that the approximations starting from the initial states (9) and (18) leading to the state (26) are extremely accurate, as long as one works within the approximation (14). For example, the fidelity of the exact state with respect to (26) is at the $99.8\%$ level choosing $\alpha=1,\beta=i$, and improves further as $|\alpha|,|\beta|$ are increased. Most of the approximation in obtaining the teleported states shown in Table 1 arise from the final measurement, to perform the collapse in (26). We therefore start with the expressions (26) and then perform numerical computations of the displacement operation and final measurement. We focus on the $\sigma_{a}=\sigma_{b}=+$ outcome and examine the performance of the teleportation protocol. For the case $\sigma_{a}^{\prime}=+$ there is no correction necessary such that $C_{+++}=I_{L}=I$. Other measurement outcomes give similar results.
@@ 4 Teleportation performance using single-mode parity measurements
×$\times C^{\dagger}_{\sigma_{a}\sigma_{b}\sigma_{a}^{\prime}}|(\mu|0_{L}\rangle_{\beta}+\nu|1_{L}\rangle_{\beta}),$ (43)
@@
is the state on mode $c$ after tracing out modes $a,b$. $C_{\sigma_{a}\sigma_{b}\sigma_{a}^{\prime}}$ is the correction operation that appears in Table 1, that is physically implemented using the operations given in Sec. 3.5. The state after applying the displacement and dispersive measurement is
@@
4.1 $\alpha\in\mathbb{R}$ and $\beta\in\mathbb{I}$, $N=1$ measurements
In Fig. 4(a) we plot the fidelities for various choices of $\alpha\in\mathbb{R}$ and $\beta\in\mathbb{I}$ for the fixed interaction time $\tau=\pi/(4\alpha^{2})$. We note that this interaction time is not the optimal time as derived in (28) unless $|\beta|=|\alpha|$. For the final dispersive measurement we consider the measurement outcome $\sigma_{a}^{\prime}=+$ corresponding to a collapse on the |±$|\pm\chi\rangle$ states (see Fig. 3(e)). We see that high fidelities are obtained for various choices of $\alpha,\beta$. The best fidelities are obtained in regions with larger $|\alpha|$ and $|\beta|$, and it is maximized when $\beta=i\alpha$. Even higher fidelities are expected with further increase of $|\alpha|$ and $|\beta|$.
4.2 $\alpha\in\mathbb{R}$ and $\beta\in\mathbb{I}$ , $N=1000$ measurements
Fig. 4(b) shows the results for the same parameters as Fig. 4(a), but the dispersive measurements at the end are repeated $N=1000$ times. We again use the fixed interaction time $\tau=\pi/(4\alpha^{2})$, which is suboptimal as derived in (28) unless $|\beta|=|\alpha|$. Here we consider the outcome $k_{+}=N,k_{-}=0$. This corresponds to the Gaussian peak at the position $x=\tau n=0$ according to (34). This collapses the state on the vacuum state, which is originally the state $|-\chi\rangle$ before the displacement, and hence suppresses the |±$|\pm\bar{\chi}\rangle$ (see Fig. 3(e)). We see that increasing the number of measurements produces an improvement of the fidelity in comparison to Fig. 4(a). As can be seen in Fig. 4(b), near unit fidelities are attainable in a larger range of parameters. The reason for this is that for the interaction time $\tau=\pi/(4\alpha^{2})$, the measurement function depends only on the parameter $\alpha$, and the larger $\alpha$ ensures that $M_{+}$ gives zero on |±$|\pm\chi\rangle$ states, thereby leading to the maximum suppression of the |±$|\pm\chi\rangle$ states.
@@ 4.2 $\alpha\in\mathbb{R}$ and $\beta\in\mathbb{I}$ , $N=1000$ measurements
4.3 Dependence on the interaction time $\tau=\pi/(\xi\bar{n})$
To verify the extent to which (28) can predict the optimal time, we calculate the fidelity for various interaction times. In Fig. 5(a) we plot the fidelity as a function of $\xi$, where $\tau=\pi/(\xi\bar{n})$. We find that the optimal points are at the point $\xi\approx 2.05$ for all the chosen $\beta\in\mathbb{R}$, and the best fidelity is at the level of 95%. This is close to the formula (28), which predicts $\xi=2$. The discrepancy arises due to (28) being based on the average photon number of the |±$|\pm\bar{\chi}\rangle$ states. Small deviations of this may produce slightly better results.
In Fig. 5(b) we show the fidelities as a function of $\xi$ for parameter choices with $\beta\in\mathbb{I}$. We find that the optimal points are sensitive to the parameter $\xi$ unlike we observed in Fig. 5(a). The reason for this is investigated in more detail in Appendix A. In short, when $\beta\in\mathbb{I}$, the effect of the peak $|\chi\rangle$ in Fig. 3(e) is closer to the |±$|\pm\bar{\chi}\rangle$ peak, and there is a stronger interference effect between the superposition of coherent states, which affects the optimal time. Nevertheless, $\tau=\pi/(2\bar{n})$ gives an excellent approximation for the optimal time, with similar fidelities being achieved as the $\beta\in\mathbb{R}$ case.
4.4 Dependence on the number of measurements $N$
We have already see in Sec. 4.2 that higher fidelities are obtained when a large number of measurements are made in the final step. Here we examine the dependence with the total number of measurements $N$. The fidelity as a function of $N$ is shown in Fig. 6. We choose the nearly optimal interaction time of $\tau=\pi/(2\bar{n})$. We see that the fidelity approaches 1 as the number of measurements is increased, as the measurements eventually collapse the state in the Fock basis and distinguish between the |±$|\pm\chi\rangle$ and |±$|\pm\bar{\chi}\rangle$ states. As shown in the insets of Fig. 6(a)(b), the larger values $|\beta|$ exhibit a faster convergence rate. Comparing the convergence for both real and imaginary $\beta$, we see that the real cases shows slightly better results. The results show that for most $\beta$, approximately $N=10$ measurements is sufficient to achieve over $99\%$ fidelity.
@@ 4.4 Dependence on the number of measurements $N$
4.5 Average over dispersive measurement outcomes
@@
5 Teleportation performance using joint parity measurements
In the previous section, we examined the performance of the teleportation using single-mode parity measurements. In this section, we examine the performance using joint parity measurements (Variant 2), as described in Sec. 3.4.2.
As with the previous section, we start with the state (26) and then perform the displacement $D(\chi)$, followed by the joint parity measurement of modes $a$ and $b$. We use the same parameters as we performed in Sec. 4 so that we can compare the results and determine which measurement scheme achieves superior fidelity. The postselected outcomes that we choose are for the case $\sigma_{ab}=+$. Here, we only present a few representative results for comparative analysis with the the single-mode parity measurement results.
Similarly to the definition in Sec. 4, we now define the fidelity for the outcome $(\sigma_{ab},\sigma_{a}^{\prime})$ with joint parity measurements:
@@ 5 Teleportation performance using joint parity measurements
×$\times C^{\dagger}_{\sigma_{ab}\sigma_{a}^{\prime}}|(\mu|0_{L}\rangle_{\beta}+\nu|1_{L}\rangle_{\beta})$ (48)
@@
5.1 $\alpha\in\mathbb{R}$ and $\beta\in\mathbb{I}$ , $N=1$ measurements
In Fig. 8(a) we plot the fidelities for various choices of $\alpha\in\mathbb{R}$ and $\beta\in\mathbb{I}$ with fixed interaction time $\tau=\pi/(4\alpha^{2})$. This is the equivalent of Fig. 4(a) but using joint parity measurements. As before, this interaction time is not the optimal time as derived in (28) unless $|\beta|=|\alpha|$. For the final dispersive measurement we consider the measurement outcome $\sigma_{a}^{\prime}=+$ corresponding to a collapse on the |±$|\pm\chi\rangle$ states (see Fig. 3(e)). We again see high fidelities are obtained for various choices of $\alpha,\beta$. The best fidelities areas occur along the line $\beta=i\alpha$, with the highest fidelities occurring for large $|\alpha|,|\beta|$. Generally, very similar results are obtained to the single-mode parity measurement case. We thus conclude that the effect on the two states (26) and (LABEL:jointparitymeasstate) are generally similar. We give the further discussion of this in Appendix B.2
@@ 5.1 $\alpha\in\mathbb{R}$ and $\beta\in\mathbb{I}$ , $N=1$ measurements
5.2 $\alpha\in\mathbb{R}$ and $\beta\in\mathbb{I}$ , $N=1000$ measurements
5.3 Dependence on the interaction time $\tau=\pi/(\xi\bar{n})$
@@
6 Conclusions

### v2

@@ III.4.2 Variant 2: Joint parity measurements
On closer inspection of Table 1, examining the outcomes for $\sigma_{a},\sigma_{b}$, one may see that there is indeed a redundancy that can be eliminated. We see that for the four mode $c$ states, the product $\sigma_{a}\sigma_{b}$ yields the same value. This suggests that by performing a joint parity measurement of $\sigma_{ab}=\sigma_{a}\sigma_{b}$, without measuring $\sigma_{a},\sigma_{b}$ individually, one can achieve the same result as Table 1. Such joint parity measurements have also been demonstrated experimentally Wang et al. (2016); Huai et al. (2024). Such a method can be calculated by applying the joint parity projection operator $P_{\sigma_{ab}}$ on the state (20).
@@
Similarly to (26), this involves a superposition of two versions of the initial state. The difference here is that modes $a$ and $b$ are left in one of the Bell states. The superposition of these terms can be further collapsed by distinguishing between the amplitudes $\chi$ and $\bar{\chi}$, using a procedure similar to that described in Sec. III.4.1. In Table 2, we summarize the various outcomes for the joint parity measurement and the displaced measurement.
@@
$\mu|0_{L}\rangle_{\beta}+\nu|1_{L}\rangle_{\beta}$ ++ ++ $I_{L}$
$\mu|0_{L}\rangle_{\beta}-\nu|1_{L}\rangle_{\beta}$ ++ - $Z_{L}$
$\mu|1_{L}\rangle_{\beta}+\nu|0_{L}\rangle_{\beta}$ - ++ $X_{L}$
$\mu|1_{L}\rangle_{\beta}+\nu|0_{L}\rangle_{\beta} -$ - $X_{L}Z_{L}$
@@ III.5 Correction operation
IV Teleportation performance using single-mode parity measurements
In this section, we numerically analyze the teleportation protocol as introduced in Sec. III. Of the two variants of the teleportation protocol involving the single-mode parity measurements (Variant 1, Sec. III.4.1) and the joint parity measurements (Variant 2, Sec. III.4.2), in this section we consider the former. The latter will be analyzed in Sec. V.
We have numerically verified that the approximations starting from the initial states (9) and (18) leading to the state (26) are extremely accurate, as long as one works within the approximation (14). For example, the fidelity of the exact state with respect to (26) is at the $99.8\%$ level choosing $\alpha=1,\beta=i$, and improves further as $|\alpha|,|\beta|$ are increased. In most systems, the amplitudes of the cat states are at the few photon level Vlastakis et al. (2013); Sun et al. (2014); Kirchmair et al. (2013); Bild et al. (2023); Yang et al. (2025), such that the values of $|\alpha|$ and $|\beta|$ that we chose are in the range $[3,4]$. As can be seen for $|\alpha|=3$ in Ref. Yang et al. (2025), the approximate relation (14) performs well experimentally. Most of the approximation in obtaining the teleported states shown in Table 1 arise from the final measurement, to perform the collapse in (26). We therefore start with the expressions (26) and then perform numerical computations of the displacement operation and final measurement. We focus on the $\sigma_{a}=\sigma_{b}=+$ outcome and examine the performance of the teleportation protocol. For the case $\sigma_{a}^{\prime}=+$ there is no correction necessary such that $C_{+++}=I_{L}=I$. Other measurement outcomes give similar results.
@@ IV Teleportation performance using single-mode parity measurements
$\times C^{\dagger}_{\sigma_{a}\sigma_{b}\sigma_{a}^{\prime}}|(\mu|0_{L}\rangle_{\beta}+\nu|1_{L}\rangle_{\beta}),$ (43)
@@
is the state on mode $c$ after tracing out modes $a,b$. $C_{\sigma_{a}\sigma_{b}\sigma_{a}^{\prime}}$ is the correction operation that appears in Table 1, that is physically implemented using the operations given in Sec. III.5. The state after applying the displacement and dispersive measurement is
@@
IV.1 $\alpha\in\mathbb{R}$ and $\beta\in\mathbb{I}$, $N=1$ measurements
In Fig. 4(a) we plot the fidelities for various choices of $\alpha\in\mathbb{R}$ and $\beta\in\mathbb{I}$ for the fixed interaction time $\tau=\pi/(4\alpha^{2})$. We note that this interaction time is not the optimal time as derived in (28) unless $|\beta|=|\alpha|$. For the final dispersive measurement we consider the measurement outcome $\sigma_{a}^{\prime}=+$ corresponding to a collapse on the $|\pm\chi\rangle$ states (see Fig. 3(e)). We see that high fidelities are obtained for various choices of $\alpha,\beta$. The best fidelities are obtained in regions with larger $|\alpha|$ and $|\beta|$, and it is maximized when $\beta=i\alpha$. Even higher fidelities are expected with further increase of $|\alpha|$ and $|\beta|$.
IV.2 $\alpha\in\mathbb{R}$ and $\beta\in\mathbb{I}$ , $N=1000$ measurements
Fig. 4(b) shows the results for the same parameters as Fig. 4(a), but the dispersive measurements at the end are repeated $N=1000$ times. We again use the fixed interaction time $\tau=\pi/(4\alpha^{2})$, which is suboptimal as derived in (28) unless $|\beta|=|\alpha|$. Here we consider the outcome $k_{+}=N,k_{-}=0$. This corresponds to the Gaussian peak at the position $x=\tau n=0$ according to (34). This collapses the state on the vacuum state, which is originally the state $|-\chi\rangle$ before the displacement, and hence suppresses the $|\pm\bar{\chi}\rangle$ (see Fig. 3(e)). We see that increasing the number of measurements produces an improvement of the fidelity in comparison to Fig. 4(a). As can be seen in Fig. 4(b), near unit fidelities are attainable in a larger range of parameters. The reason for this is that for the interaction time $\tau=\pi/(4\alpha^{2})$, the measurement function depends only on the parameter $\alpha$, and the larger $\alpha$ ensures that $M_{+}$ gives zero on $|\pm\chi\rangle$ states, thereby leading to the maximum suppression of the $|\pm\chi\rangle$ states.
@@ IV.2 $\alpha\in\mathbb{R}$ and $\beta\in\mathbb{I}$ , $N=1000$ measurements
IV.3 Dependence on the interaction time $\tau=\pi/(\xi\bar{n})$
To verify the extent to which (28) can predict the optimal time, we calculate the fidelity for various interaction times. In Fig. 5(a) we plot the fidelity as a function of $\xi$, where $\tau=\pi/(\xi\bar{n})$. We find that the optimal points are at the point $\xi\approx 2.05$ for all the chosen $\beta\in\mathbb{R}$, and the best fidelity is at the level of 95%. This is close to the formula (28), which predicts $\xi=2$. The discrepancy arises due to (28) being based on the average boson number of the $|\pm\bar{\chi}\rangle$ states. Small deviations of this may produce slightly better results.
In Fig. 5(b) we show the fidelities as a function of $\xi$ for parameter choices with $\beta\in\mathbb{I}$. We find that the optimal points are sensitive to the parameter $\xi$ unlike we observed in Fig. 5(a). The reason for this is investigated in more detail in Appendix A. In short, when $\beta\in\mathbb{I}$, the effect of the peak $|\chi\rangle$ in Fig. 3(e) is closer to the $|\pm\bar{\chi}\rangle$ peak, and there is a stronger interference effect between the superposition of coherent states, which affects the optimal time. Nevertheless, $\tau=\pi/(2\bar{n})$ gives an excellent approximation for the optimal time, with similar fidelities being achieved as the $\beta\in\mathbb{R}$ case.
IV.4 Dependence on the number of measurements NN
We have already see in Sec. IV.2 that higher fidelities are obtained when a large number of measurements are made in the final step. Here we examine the dependence with the total number of measurements $N$. The fidelity as a function of $N$ is shown in Fig. 6. We choose the nearly optimal interaction time of $\tau=\pi/(2\bar{n})$. We see that the fidelity approaches 1 as the number of measurements is increased, as the measurements eventually collapse the state in the Fock basis and distinguish between the $|\pm\chi\rangle$ and $|\pm\bar{\chi}\rangle$ states. As shown in the insets of Fig. 6(a)(b), the larger values $|\beta|$ exhibit a faster convergence rate. Comparing the convergence for both real and imaginary $\beta$, we see that the real cases shows slightly better results. The results show that for most $\beta$, approximately $N=10$ measurements is sufficient to achieve over $99\%$ fidelity.
@@ IV.4 Dependence on the number of measurements NN
IV.5 Average over dispersive measurement outcomes
@@
V Teleportation performance using joint parity measurements
In the previous section, we examined the performance of the teleportation using single-mode parity measurements. In this section, we examine the performance using joint parity measurements (Variant 2), as described in Sec. III.4.2.
As with the previous section, we start with the state (26) and then perform the displacement $D(\chi)$, followed by the joint parity measurement of modes $a$ and $b$. We use the same parameters as we performed in Sec. IV so that we can compare the results and determine which measurement scheme achieves superior fidelity. The postselected outcomes that we choose are for the case $\sigma_{ab}=+$. Here, we only present a few representative results for comparative analysis with the the single-mode parity measurement results.
Similarly to the definition in Sec. IV, we now define the fidelity for the outcome $(\sigma_{ab},\sigma_{a}^{\prime})$ with joint parity measurements:
@@ V Teleportation performance using joint parity measurements
$\times C^{\dagger}_{\sigma_{ab}\sigma_{a}^{\prime}}|(\mu|0_{L}\rangle_{\beta}+\nu|1_{L}\rangle_{\beta})$ (48)
@@
V.1 $\alpha\in\mathbb{R}$ and $\beta\in\mathbb{I}$ , $N=1$ measurements
In Fig. 8(a) we plot the fidelities for various choices of $\alpha\in\mathbb{R}$ and $\beta\in\mathbb{I}$ with fixed interaction time $\tau=\pi/(4\alpha^{2})$. This is the equivalent of Fig. 4(a) but using joint parity measurements. As before, this interaction time is not the optimal time as derived in (28) unless $|\beta|=|\alpha|$. For the final dispersive measurement we consider the measurement outcome $\sigma_{a}^{\prime}=+$ corresponding to a collapse on the $|\pm\chi\rangle$ states (see Fig. 3(e)). We again see high fidelities are obtained for various choices of $\alpha,\beta$. The best fidelities areas occur along the line $\beta=i\alpha$, with the highest fidelities occurring for large $|\alpha|,|\beta|$. Generally, very similar results are obtained to the single-mode parity measurement case. We thus conclude that the effect on the two states (26) and (LABEL:jointparitymeasstate) are generally similar. We give the further discussion of this in Appendix B.2
V.2 $\alpha\in\mathbb{R}$ and $\beta\in\mathbb{I}$ , $N=1000$ measurements
@@
V.3 Dependence on the interaction time $\tau=\pi/(\xi\bar{n})$
@@
VI Decoherence
The study of decoherence is crucial for the realization of quantum teleportation, as any experiment is inevitably affected by the environment. In particular, for teleportation protocols based on Schrödinger cat states, their macroscopic nature make them more sensitive to decoherence compared to other types of quantum states Zurek (2003). In this section, we introduce decoherence into our protocol, and evaluate its performance to confirm that the protocol remains practical within experimentally achievable limits.
Here, we consider the loss channel, which can be characterized by the a set of Kraus operators Nielsen and Chuang (2010)
$A_{k}=\sum_{n=k}^{\infty}\sqrt{n\choose k}\sqrt{\gamma}^{n-k}\sqrt{1-\gamma}^{k}|n-k\rangle\langle k|,$ (52)
where $1-\gamma$ is the loss probability, $k$ is the number of lost bosons. Suppose that the loss occurs on mode $a$, affecting the ability to distinguish between $\chi$ and $\bar{\chi}$ effectively. Note that we only discuss the case of Variant 1, as the results obtained for Variant 2 are similar. We apply $A_{k}$ to (26) on mode $a$ such that the state is
$|\psi^{k}_{\sigma_{a}\sigma_{b}}\rangle=\frac{A_{k}|\psi_{\sigma_{a}\sigma_{b}}\rangle}{\sqrt{p_{k}}\sqrt{p_{\sigma_{a}\sigma_{b}}}},$ (53)
where
$p_{k}=\langle\psi_{\sigma_{a}\sigma_{b}}|A^{\dagger}_{k}A_{k}|\psi_{\sigma_{a}\sigma_{b}}\rangle.$ (54)
is the probability that $k$ bosons are lost. We now define the fidelity for the measurement outcome $(\sigma_{a},\sigma_{b},\sigma_{a}^{\prime})$ as
$F^{k}_{\sigma_{a}\sigma_{b}\sigma_{a}^{\prime}}=$ $(\mu^{*}\langle 0_{L}|_{\beta}+\nu^{*}\langle 1_{L}|)C_{\sigma_{a}\sigma_{b}\sigma_{a}^{\prime}}\rho^{k}_{\sigma_{a}^{\prime}|\sigma_{a}\sigma_{b}}$
$\times C^{\dagger}_{\sigma_{a}\sigma_{b}\sigma_{a}^{\prime}}|(\mu|0_{L}\rangle_{\beta}+\nu|1_{L}\rangle_{\beta}),$ (55)
where
$\rho^{k}_{\sigma_{a}^{\prime}|\sigma_{a}\sigma_{b}}=\frac{\text{Tr}_{a,b}|\psi^{k}_{\sigma_{a}^{\prime}|\sigma_{a}\sigma_{b}}\rangle\langle\psi^{k}_{\sigma_{a}^{\prime}|\sigma_{a}\sigma_{b}}|}{p^{k}_{\sigma_{a}^{\prime}|\sigma_{a}\sigma_{b}}}$ (56)
is the state on mode $c$ after tracing out modes $a$ and $b$. $C_{\sigma_{a}\sigma_{b}\sigma_{a}^{\prime}}$ is the same correction operation as we introduced in Sec. IV and V. The state after applying the displacement and dispersive measurement is
$|\psi^{k}_{\sigma_{a}^{\prime}|\sigma_{a}\sigma_{b}}\rangle=M_{\sigma_{a}^{\prime}}D(\chi)|\psi^{k}_{\sigma_{a}\sigma_{b}}\rangle,$ (57)
and the conditional probability is
$p^{k}_{\sigma_{a}^{\prime}|\sigma_{a}\sigma_{b}}=\langle\psi^{k}_{\sigma_{a}^{\prime}|\sigma_{a}\sigma_{b}}|\psi^{k}_{\sigma_{a}^{\prime}|\sigma_{a}\sigma_{b}}\rangle.$ (58)
We then obtain the fidelity averaged over $k$
$F^{\text{av}}_{\sigma_{a}\sigma_{b}\sigma_{a}^{\prime}}=\sum_{k}p_{k}F_{\sigma_{a}\sigma_{b}\sigma_{a}^{\prime}}^{k}.$ (59)
Figure 10: Fidelity (59) as a function of $1-\gamma$ for Variant 1 teleportation with single-mode parity measurements. Parameters are chosen such that (a) $\beta\in\mathbb{R},N=1,k_{+}=1,k_{-}=0$; (b) $\beta\in\mathbb{I},N=1,k_{+}=1,k_{-}=0$; (c) $\beta\in\mathbb{R},N=1000,k_{+}=1000,k_{-}=0$; (d) $\beta\in\mathbb{I},N=1000,k_{+}=1000,k_{-}=0$. Common parameters are $\mu=1/2,\nu=\sqrt{3}/2,\alpha=4,\tau=\pi/(2\bar{n})$.
Fig. 10 shows the average fidelity (59) versus the probability of loss $1-\gamma$ under single-mode measurement for various parameters. First, we see that for loss probabilities in the range of a few $\sim\%$, the fidelities are largely unchanged. For larger loss probabilities, we observe that the larger $|\beta|$ of cat states have a faster decoherence rate, which is expected as larger cat states are more susceptible to loss. We see that the fidelity converges with loss rate for $\beta\in\mathbb{R}$ (Fig. 10(a)). This is mainly because the ability of distinguishing the distribution no longer changes significantly. However, for $\beta\in\mathbb{I}$ the fidelity seems to converge at a higher loss rate (Fig. 10(b)). For example, the fidelity within the choice of parameters $\beta\in\mathbb{I}$ is significantly higher than that using $\beta\in\mathbb{R}$ for the plotted range. Hence, we conclude that the protocol is more robust in this case.
In previous sections, we showed that by using repeated dispersive measurements, one can obtain a fidelity approaching 1. Similarly, we now examine the multiple measurements case in the presence of loss. The results are shown in Fig. 10(c)(d). We find that compared to single measurements, the fidelity is substantially improved, especially for larger choices of the amplitude $|\beta|$. For example, when $|\beta|=4$, the system can almost completely resist loss up to probabilities $1-\gamma<\sim 15\%$, as illustrated in Fig. 10(c). Likewise, Fig. 10(d) shows that choosing $\beta\in\mathbb{I}$ leads to significantly more robust results. Choosing $\beta=4i$ and performing multiple measurements can yield a nearly perfect fidelity, even the probability of loss is up to 35%.
VII Conclusions